OpenStudy (anonymous):

Evaluate. a^0 + 3^2 / 2^4 - b^0

OpenStudy (anonymous):

Still here?

OpenStudy (anonymous):

a^0 = 1 3^2 = 9 1 + 9 = 10 2^4 = 16 b^0 = 1 16-1=15 answer: 10/15

OpenStudy (anonymous):

It says in my book that the answer is 4 though :(

OpenStudy (anonymous):

You should ask your teacher about that; I went over it twice as a second opinion, and 10/15 is correct, though you can simplify it to 2/3.

OpenStudy (anonymous):

Maybe the book was wrong :( Thank you though! :)

OpenStudy (anonymous):

Hmm im pretty sure the book is wrong. (Which happens. My math book is full of mistakes in the back of the book)

OpenStudy (anonymous):

sorry for the super late response.

OpenStudy (anonymous):

Odds are that the book is indeed wrong. The other explanation is that you've typed the question up incorrectly. Double check that, and if you've typed it up right, then the book is almost certainly wrong on this one.

OpenStudy (anonymous):

^ that is the only other option.

OpenStudy (anonymous):

Oh shoot it's suppose to be 3^3 instead of 3^2 D: sorry!!

OpenStudy (anonymous):

Ah. In that case, the book answer is STILL wrong, because we have 1 + 27 / 16 - 1 or 28/15.

OpenStudy (anonymous):

haha yup. 28/15. Are you looking at the right spot for the answer?

OpenStudy (anonymous):

Omg, im so sorry again! The aswer on the book is 2/3

OpenStudy (anonymous):

Then the original answer we gave is correct. :)

OpenStudy (anonymous):

10/15 simplified to 2/3

OpenStudy (anonymous):

THANK YOU SO MUCH! :) Im sorry for all the mistakes

OpenStudy (anonymous):

Not a problem :) Now you know what to check if this situation arises again.